OpenStudy (anonymous):

The lateral area of a cube is two hundred fifty-six square kilometers. What is its side length? 4 kilometers 8 kilometers 24 kilometers 32 kilometers

rishavraj (rishavraj):

see the lateral surface area of cube = 4 a^2 = 256...get a

OpenStudy (anonymous):

i don't know what lateral area is.

TheSmartOne (thesmartone):

Lateral area of a cube is 4a^2 equate it to equal 256 so 4a^2=256 divide by 4 on both sides a^2=64 a=8

OpenStudy (anonymous):

thank you!!!

TheSmartOne (thesmartone):

np